# 70% Size, 100% Accuracy: Lossless LLM Compression for Efficient GPU Inference via Dynamic-Length Float

### v1

@@ Abstract
Large Language Models (LLMs) have grown rapidly in size, creating significant challenges for efficient deployment on resource-constrained hardware. In this paper, we introduce Dynamic-Length Float (DFloat11), a lossless compression framework that reduces LLM size by 30% while preserving outputs that are bit-for-bit identical to the original model. DFloat11 is motivated by the low entropy in the BFloat16 weight representation of LLMs, which reveals significant inefficiency in existing storage format. By applying entropy coding, DFloat11 assigns dynamic-length encodings to weights based on frequency, achieving near information-optimal compression without any loss of precision. To facilitate efficient inference with dynamic-length encodings, we develop a custom GPU kernel for fast online decompression. Our design incorporates the following: (i) decomposition of memory-intensive lookup tables (LUTs) into compact LUTs that fit in GPU SRAM, (ii) a two-phase kernel for coordinating thread read/write positions using lightweight auxiliary variables, and (iii) transformer-block-level decompression to minimize latency. Experiments on recent models, including Llama-3.1, Qwen-2.5, and Gemma-3, validates our hypothesis that DFloat11 achieves around 30% model size reduction while preserving bit-for-bit exact outputs. Compared to a potential alternative of offloading parts of an uncompressed model to the CPU to meet memory constraints, DFloat11 achieves 1.9–38.8$\times$ higher throughput in token generation. With a fixed GPU memory budget, DFloat11 enables 5.3–13.17$\times$ longer context lengths than uncompressed models. Notably, our method enables lossless inference of Llama-3.1-405B, an 810GB model, on a single node equipped with 8$\times$80GB GPUs. Our code and models are available at https://github.com/LeanModels/DFloat11.
@@ 1 Introduction
Table 1: Comparison of running LLMs on specific GPUs using BFloat16 and our proposed DFloat11 format. Note that DFloat11 produces outputs that are 100% identical to those of the original BFloat16 models.
Model GPU Type Method Successfully Run? Required Memory
8$\times$×H100-80G BFloat16 811.71 GB
Llama-3.1-405B-Instruct DFloat11 (Ours) 551.22 GB
Llama-3.3-70B-Instruct 1$\times$×H200-141G BFloat16 141.11 GB
DFloat11 (Ours) 96.14 GB
QwQ-32B 1$\times$×A6000-48G BFloat16 65.53 GB
DFloat11 (Ours) 45.53 GB
Qwen2.5-32B-Instruct 1$\times$×A6000-48G BFloat16 65.53 GB
DFloat11 (Ours) 45.53 GB
DeepSeek-R1-Distill-Llama-8B 1$\times$×RTX 5080-16G BFloat16 16.06 GB
DFloat11 (Ours) 11.23 GB
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language processing (NLP) tasks [38]. However, their rapidly increasing sizes create substantial obstacles to efficient deployment and inference, especially in environments with limited computational or memory resources. For example, Llama-3.1-405B [11] has 405 billion parameters in 16-bit Brain Float (BFloat16) format and requires about 810 GB of memory for full inference, exceeding the capacity of a typical high-end GPU server (e.g., DGX A100/H100 with 8$\times$80GB GPUs). As a result, deploying this model requires multiple nodes, making it expensive and inaccessible. In this work, we present a solution that compresses any BFloat16 model to approximately 70% of its original size while preserving 100% of its accuracy on any task.
1.1 Lossy Compression Degrades Output Quality, and Lossless Compression Lacks GPU Efficiency
To address the growing model size of LLMs, quantization techniques [7, 21] are commonly employed, converting high-precision weights into lower-bit representations. This significantly reduces memory footprint and computational requirements, facilitating faster inference and deployment in resource-constrained environments. However, quantization is inherently a lossy compression technique, introducing a fundamental drawback: it inevitably alters the output distribution of LLMs, thus impacting model accuracy and reliability.
In contrast, lossless compression techniques preserve the exact original weights of large-scale LLMs while effectively reducing their size, ensuring the model’s output distribution remains identical to that of the uncompressed representation (e.g., BFloat16). However, existing lossless methods primarily focus on improving storage efficiency for LLMs, such as shrinking model checkpoints [13, 16] or optimizing performance for specialized hardware like FPGAs [40]. While these methods indeed benefit scenarios like efficient checkpoint rollbacks during training [32] or accelerated downloads from model repositories such as HuggingFace [16], their advantages typically do not extend effectively to general-purpose GPU-based LLM inference.
1.2 Dynamic-Length Float (DFloat): A Data Format for Lossless GPU Inference
Our key insight is that the BFloat16 (BF16) format used to store LLM weights is information-inefficient. Through an analysis of the frequency distribution of BFloat16 components (sign, exponent, and mantissa) in recent LLMs, we leverage the observation that BFloat16 exponents carry significantly less information than their allocated bit width, with an entropy of approximately 2.6 bits compared to the assigned 8 bits [13]. This highlights an opportunity for substantial lossless compression. Leveraging entropy coding techniques such as Huffman coding [17], which assign shorter codes to more frequent symbols, we can achieve roughly 30% compression without any loss in information. However, such a compressed representation introduces considerable challenges for efficient GPU inference.
Although theoretically appealing, efficient inference with entropy-coded weights on GPUs presents significant challenges. During inference, losslessly compressed weight matrices must be decompressed on-the-fly back into their original BFloat16 format for matrix multiplications, and this decompression step introduces a critical performance bottleneck. Traditional Huffman decoding algorithms rely on sequential, bit-by-bit traversal of a Huffman tree, making them inherently unsuitable for GPU’s massively parallel architecture. Assigning a single GPU thread for decompression, however, results in severe under-utilization of GPU resources and increased latency. Existing entropy coding-based compression methods, ranging from established CNN-focused approaches [13] to recent adaptations for LLMs [16, 14, 40], have yet to deliver significant inference efficiency improvements in GPU-based scenarios relevant to most end-users.
To enable efficient inference with entropy-coded weights on GPUs, we introduce a novel data representation called Dynamic-Length Float (DFloat), along with a GPU kernel designed for fast, online decompression of 11-bit DFloat weights (DFloat11) in a massively parallel manner. The DFloat11 decompression kernel comprises three core components: ➊ Efficient decoding of entropy-coded weights using compact lookup tables (LUTs) stored in GPU shared memory (SRAM). A single, monolitic LUT for decoding 32-bit Huffman codes would require roughly 4.29 billion entries (as discussed in Section 3.3.1), making it prohibitively memory-intensive. To mitigate this, we decompose it into multiple compact LUTs that fit within GPU SRAM to enable fast access. ➋ Precise and efficient identification of “READ” positions in encoded weights and “WRITE” positions for decoded weights. Since entropy-encoded weights have variable bit widths and are tightly packed, determining correct read/write offsets for each thread is challenging. We resolve this through a two-phase kernel design, which employs a minimal set of auxiliary variables to efficiently coordinate thread-specific input and output positions. ➌ Performing matrix decompression in batch for improved GPU resource utilization. Decompressing individual weight matrices leads to poor GPU utilization due to their relatively small size. To address this, we decompress weights at the transformer-block level, significantly enhancing throughput and reducing inference latency.
We summarize our contributions as follows:
@@ 1.2 Dynamic-Length Float (DFloat): A Data Format for Lossless GPU Inference
  • We propose Dynamic-Length Float (DFloat11), a data format that losslessly compresses BFloat16 weights of LLMs down to approximately 11 bits. By leveraging the information inefficiency of the BFloat16 representation, we achieve roughly a 30% “free” model-size reduction while preserving bit-for-bit identical outputs.
@@
  • We introduce optimized algorithmic designs to enable efficient GPU inference with DFloat11-compressed models. By carefully exploiting GPU memory and computational hierarchies, we develop hardware-aware algorithms for efficient online inference of DFloat11-compressed models.
  • We extensively evaluate our method on popular LLMs and large reasoning models, including Llama-3.1, Qwen2.5, QwQ-32B, Mistral, Gemma-2, Gemma-3, DeepSeek-R1-Distill-Qwen, and DeepSeek-R1-Distill-Llama [11, 36, 31, 30, 27, 26, 12]. Experimental results demonstrate that our method consistently achieves about 30% compression without altering outputs. For example, our approach reduces the hardware requirements for running Llama-3.1-405B from two GPU nodes down to a single node, equipped with 8$\times$80GB NVIDIA A100 GPUs, without any loss in accuracy or changes in output distribution.
2 Motivation: Is Lossless Compression of LLMs Worth Studying?

### v2

@@ Abstract
[…]
@@ 1 Introduction
Foundation models, such as Large Language Models (LLMs) and Diffusion Models (DMs), have demonstrated remarkable capabilities across a wide range of Natural Language Processing (NLP) yang2024harness_survey and Computer Vision (CV) tasks yang2023diffusion . However, their huge model sizes create substantial obstacles for efficient deployment, especially in memory-constrained environments. For example, a competitive recent LLM, Llama 3.1 405B grattafiori2024llama , has 405 billion parameters in 16-bit Brain Float (BFloat16) format and requires about 810 GB of memory for full inference, exceeding the capacity of a typical high-end GPU server (e.g., DGX A100/H100 with $\times8$0GB GPUs). As a result, deploying this model requires multiple nodes, making it expensive and inaccessible. In this work, we present a solution that compresses any BFloat16 model to approximately 70% of its original size while preserving 100% of its accuracy on any task.
Model compression via quantization has limitations. Quantization is a type of lossy compression method that lowers the precision of model weights by converting them into lower bit-width representations frantar2022gptq ; lin2024awq ; li2023q ; sui2024bitsfusion . Although it can significantly reduce memory usage and often improve inference speed, quantization is not a one-size-fits-all solution and presents several key limitations: ➊ Accuracy degradation. By design, quantization introduces approximation errors. The degree of accuracy loss depends on multiple factors, including the base model, quantization method, evaluation benchmark, and target bit-width eval_quantized . These interactions make it difficult to predict or quantify the impact comprehensively. Even mild quantization can noticeably degrade performance. For example, applying 8-bit SmoothQuant xiao2023smoothquant to DeepSeek-R1-Distill-Qwen-1.5B guo2025deepseek results in a 9.09% drop in average accuracy across reasoning tasks liu2025quant_reason_bench . ➋ Behavioral shifts. Even when overall accuracy metrics appear roughly unchanged, quantized models may behave differently from their full-precision counterparts. For instance, Dutta et al. dutta2024accuracy_not observe a phenomenon called flips, where quantized models produce answers that change from correct to incorrect and vice versa. This indicates that quantization can significantly alter model behavior, even when standard accuracy metrics show minimal change. For example, the W8A16 GPTQ-quantized Qwen2-1.5Bfrantar2022gptq ; yang2024qwen2 exhibits only a 0.3% drop in GSM8K (8-shot) accuracy cobbe2021gsm8k , yet 6.37% of its answers flip in correctness dutta2024accuracy_not . ➌ Compliance and reliability concerns. In domains like finance or healthcare, quantized models may not satisfy regulatory or reliability standards, as their outputs may differ from those of the original models kharinaev2025invest_quant_safety . We refer readers to Appendix A for a more detailed discussion on quantization.
Existing lossless model compression does not support efficient GPU inference. Unlike lossy compression, lossless compression reduces model size while preserving the full precision of the original weights. This ensures the model’s output distribution remains identical to that of the uncompressed counterpart. However, most existing lossless methods focus on storage efficiency, such as compressing model checkpoints han2015deep_comp ; hershcovitch2024zipnn , or target specialized hardware like FPGAs yubeaton2025huffllm , rather than accelerating inference on general-purpose GPUs. While useful for tasks like checkpoint rollback during large-scale training wang2023gemini_ckpt or reducing download time from model hubs hershcovitch2024zipnn , these methods offer little to no benefit for GPU-based inference.
Our proposal, Dynamic-Length Float (DFloat11), is a lossless compression framework optimized for efficient GPU inference. We identify a key inefficiency in the commonly used BFloat16 format: its 8-bit exponent field carries only about 2.6 bits of actual information. This redundancy is consistent across a wide range of LLMs, as shown in Section 2.2. To exploit it, we apply Huffman coding huffman_coding to the exponent bits of BFloat16 weights, while leaving the sign and mantissa bits uncompressed. The resulting exponents have dynamic-length encodings: frequent values are assigned shorter codes, while rarer ones use longer codes. However, standard Huffman decoding relies on sequential bit-by-bit tree traversal, which is inefficient on GPUs due to limited parallelism. Assigning one GPU thread per decompression task leads to severe hardware underutilization and high latency. To overcome this, we design a hardware-aware algorithm that enables efficient online decompression of dynamic-length floats on GPUs. Our solution includes three key components: 1. compact, hierarchical lookup tables (LUTs) that fit in GPU SRAM to support fast, table-based Huffman decoding, 2. a two-phase GPU kernel that uses lightweight auxiliary variables to coordinate thread-level read and write operations, and 3. batched decompression at the transformer-block level to maximize throughput. We summarize our contributions as follows:
  1. 1.
@@
    We propose Dynamic-Length Float (DFloat11), a losslessly compressed floating-point format that reduces BFloat16 weights to approximately 11 bits. This yields around 30% model size reduction with bit-for-bit identical outputs.
  2. 2.
@@
    We develop optimized, hardware-aware algorithms for efficient GPU inference with DFloat11-compressed models by leveraging GPU memory and compute hierarchies.
  3. 3.
    We evaluate DFloat11 across popular LLMs and diffusion transformers, including Llama 3, Qwen 3, Mistral 3, DeepSeek R1 Distilled, FLUX.1, and Stable Diffusion 3.5 grattafiori2024llama ; qwen3 ; mistral2025small3 ; guo2025deepseek ; flux ; stablediffusion35 . Our method consistently achieves 30% compression without altering original outputs at all. Notably, it enables running Llama-3.1-405B on a single node ($\times8$0GB A100 GPUs), reducing hardware requirements by half without accuracy loss.
2 Method
In this section, we introduce our proposed floating-point format, Dynamic-Length Float (DFloat11), along with its custom decompression kernel designed for efficient GPU inference.
2.1 Preliminary
Brain Float (BFloat16) Recent state-of-the-art LLMs predominantly employ the 16-bit Brain Float format (BFloat16 or BF16) for storing weights, due to its balance of numerical precision with memory efficiency. BF16 allocates its 16 bits as follows: 1 sign bit, 8 exponent bits, and 7 mantissa bits. The numerical value represented by a BF16 number is computed as:
where $\mathrm{mantissa}$ is interpreted as a binary fractional value.
Entropy Coding Entropy coding is a core technique in lossless data compression that leverages statistical redundancy to reduce data size. Several widely used methods fall under this category, including Huffman coding huffman_coding , arithmetic coding arithmetic_coding , and Asymmetric Numeral Systems (ANS) ans . Among these, Huffman coding is one of the most widely adopted, which uses variable-length encoding to minimize the size of encoded data. It assigns shorter binary codes to more frequent symbols and longer codes to less frequent ones. The codes are decoded using a prefix-free binary tree, known as a Huffman tree. Due to the prefix-free property of Huffman codes, no code is a prefix of any other, which ensures unique decodability of the encoded bitstream without the need for delimiters. The tree is constructed based on symbol frequencies and is provably optimal for any given frequency distribution. However, decoding Huffman codes in a massively parallel manner is challenging due to its inherently sequential nature.
GPU Computation and Memory Paradigm GPUs are designed to perform computations in a massively parallel manner. A modern GPU consists of thousands of threads, which are organized into blocks and executed on streaming multiprocessors (SMs). Each block has access to a small, fast, on-chip memory called shared memory (often referred to as SRAM), which provides much lower latency and higher bandwidth than the off-chip global memory, commonly known as high-bandwidth memory (HBM). The capacity of shared memory is limited, typically having up to 100 KB per block. In this work, we leverage the fast access characteristics of SRAM to enable efficient on-the-fly decompression of compressed weights during inference.
Figure 1: (Left) The allocation of bits for the components of BFloat16. (Right 3) The Shannon entropy of the components (sign, exponent, mantissa) of BFloat16 weights in various LLMs.
2.2 Motivation: BFloat16 Representation is Information Inefficient
To motivate the lossless compression of LLM weights, we analyze the compressibility of the BFloat16 weights of recent LLMs. Specifically, we use Shannon entropy to quantify the information content of BFloat16 components (sign, exponent, and mantissa) for all linear projection matrices of an LLM. The Shannon entropy $H(\cdot)$ is defined as:
$H(X)=-\sum_{x\in\mathcal{X}}p(x)\log_{2}p(x)$ (2)
where $X$ is a discrete random variable with support $\mathcal{X}$, and $p:\mathcal{X}\to[0,1]$ denotes its probability mass function. We present the computed entropy values in Figure 1. As shown, the entropy of the sign and mantissa bits is close to their respective bit widths, indicating limited potential for compression. In contrast, the exponent exhibits significantly lower entropy, approximately 2.6 bits versus its allocated 8 bits, suggesting substantial opportunities for lossless compression.
To understand this discrepancy, we visualize the frequency distribution of all BFloat16 components in Figure 8 and the ranked frequency of exponent values in Figure 9, both in the Appendix. The sign and mantissa values are relatively uniform across their ranges, but the exponent distribution is highly imbalanced: only about 40 of the 256 possible 8-bit values are used, with the rest never appearing. Ranked frequencies also decay rapidly. These observations reveal the low entropy of the exponent and its potential for compression.
Figure 2: Our proposed format Dynamic-Length Float for compressing BFloat16 weights of LLMs losslessly down to 11 bits. The exponents are compressed via Huffman coding, while the sign and mantissa bits remain uncompressed.
2.3 Dynamic-Length Float: Lossless LLM Compression for Efficient GPU Inference
To address the substantial information inefficiency in the BFloat16 representation of LLM weights, we propose a lossless compression framework that encodes floating-point parameters using entropy coding. Specifically, we build a Huffman tree based on the distribution of exponents in model weights. We then compress the exponents using Huffman coding, while preserving the original signs and mantissas. Exponents are encoded and tightly bit-packed into a byte array, $\mathsf{EncodedExponent}$, while the sign and mantissa are left uncompressed and stored in a separate byte array $\mathsf{PackedSignMantissa}$. Figure 2 illustrates Dynamic-Length Float (DFloat11 or DF11), our proposed format for compactly representing BFloat16 model parameters.
The Core Challenge: Efficient GPU Inference with Compressed Weights While DFloat11 enables lossless compression of LLM weights, efficient GPU inference remains a key challenge. Entropy-coded weights use variable-length encoding and cannot be directly used in matrix multiplications. As a result, each weight matrix must be decompressed on-the-fly to its original BFloat16 format before matrix multiplication, then discarded immediately after use to conserve memory. However, traditional Huffman decoding is inherently sequential, requiring bit-by-bit tree traversal for each element, which is ill-suited for GPUs’ parallel architecture. Naively assigning a single thread for decompression leads to poor utilization and high latency. Addressing this bottleneck is essential for practical compressed inference.
In the following paragraphs, we present our solution in detail: a set of hardware-aware algorithmic designs tailored for low-latency decoding of entropy-coded weights in a massively parallel manner. Our approach consists of three key components: ➊ leveraging compact lookup tables that fit within GPU SRAM for efficient, lookup-based decoding, ➋ introducing a two-phase kernel design to coordinate read/write operations for all threads using lightweight auxiliary variables, and ➌ performing decompression at the transformer block level to minimize latency.
Figure 3: (Left) The Huffman tree is decomposed into a set of non-overlapping subtrees, each corresponding to a compact lookup table (LUT). These hierarchical LUTs reside in GPU SRAM to enable efficient Huffman decoding via array lookups. (Right) Each thread decodes $n$ bytes of encoded exponents. The array Gaps stores the bit offset of the first element assigned to each thread, while the array Block Output Positions stores the index of the first element for each thread block.
2.3.1 Efficient Decoding with Hierarchical Lookup Tables
The traditional approach to decoding Huffman codes involves reading the encoded bitstream bit by bit and traversing the Huffman tree accordingly. However, this method is inefficient on GPUs due to frequent branching and thread divergence. To enable efficient decoding on GPUs, we adopt a lookup-table-based approach huffman_gap_arrays .
Assume the maximum Huffman code length is $L$, and we construct a lookup table $\mathsf{LUT}$ of size $2^{L}$. At each index $i$, $\mathsf{LUT}$ stores the decoded exponent whose Huffman code matches the prefix of the binary representation of $i$. To decode the next exponent, we read the next $L$ bits from the encoded bitstream, interpret them as an index into $\mathsf{LUT}$, and retrieve the corresponding value. To determine how many bits to advance in the stream, we use a secondary lookup table $\mathsf{CodeLengths}$, which maps each exponent to the length of its Huffman code. A detailed example of this decoding process is provided in Section I of the Appendix.
In practice, the value of $L$ can be large. For LLMs, $L$ typically ranges from 24 to 32, resulting in a $\mathsf{LUT}$ with up to $2^{32}$ entries, which cannot fit within GPU SRAM for fast lookups. To address this, we decompose the monolithic $\mathsf{LUT}$ into a hierarchy of compact lookup tables huffman_gap_arrays . We first partition the Huffman tree into non-overlapping subtrees of height 8. Each subtree corresponds to a compact LUT that decodes 8 bits, requiring only $2^{8}=256$ entries.
Figure 3 shows an example of how a Huffman tree of height 4 can be decomposed into a hierarchy of compact LUTs, each with 4 entries. Because the LUTs are organized hierarchically, some entries must serve as references to other LUTs lower in the hierarchy. We take advantage of the sparsity in 8-bit exponent usage: although 256 values are available, typically only around 40 are used in LLMs (see Figure 9 in the Appendix). We repurpose unused values (specifically, the range 240 to 255) as pointers to other LUTs. These values correspond to extremely large magnitudes ($\pm 2^{113}$ to $\pm 2^{128}$) that do not occur in LLM weights, making them safe for use as internal markers.
We use $k$ to denote the number of compact LUTs. In our experiments, we observe that $k$ ranges from 4 to 8 for the Huffman trees built from BFloat16 exponent values. Combined with $\mathsf{CodeLengths}$, these LUTs occupy at most $(8+1)\times 256$ bytes of memory, which easily fits within SRAM and allows for fast repeated lookups.
2.3.2 Two-Phase Kernel and Lightweight Auxiliary Variables
To leverage the parallel processing capabilities of GPUs, we assign each thread to a contiguous, non-overlapping block of encoded exponents consisting of $n$ bytes ($n=8$ in our experiments). Each thread decodes elements whose Huffman codes begin within its assigned block. Since Huffman codes are variable-length, a thread may need to skip some bits at the start before decoding the first element. Similarly, the last element may span beyond the assigned byte range.
This approach introduces two key challenges: 1. The starting bit position for each thread is unclear due to the variable-length nature of Huffman codes. 2. Except for the first thread, the index of decoded elements is unknown, making it difficult to determine their correct output locations.
To address the first issue, we use a gap array huffman_gap_arrays to specify the starting bit offset for each thread. The array $\mathsf{Gaps}$ has one entry per thread, where each entry indicates the offset of the first valid Huffman code relative to the thread’s assigned starting byte. With a maximum code length of 32 bits, each offset lies in $[0,31]$ and is stored using only 5 bits.
For the second issue, maintaining an output position for each thread is straightforward but memory-intensive. Each position requires a 32-bit integer, and with tens of thousands of threads per weight matrix, this leads to significant overhead, undermining DFloat11’s compression benefits. To reduce this overhead, we store the output position only for the first element of each thread block rather than for every thread. Since each block typically contains hundreds to thousands of threads, this optimization reduces the overhead from one 32-bit integer per thread to one per block, making the memory cost negligible. Figure 3 illustrates how the gap and block-level output position arrays encode the metadata associated with the encoded exponents.
To support this design, we implement a two-phase kernel. In the first phase, each thread decodes its assigned block and counts the number of elements, without writing to the HBM. Afterward, threads within a block synchronize to compute per-thread output positions via a prefix sum over the element counts. We use the Blelloch algorithm blelloch for this step. In the second phase, each thread re-decodes the same block, this time writing decoded values to a write buffer in SRAM at the calculated positions. To avoid redundant global memory access, the encoded exponents are loaded into SRAM before the first pass. Once all decoded exponents are written to SRAM, a single batch of coalesced writes is issued to HBM. Pseudocode for the two-phase kernel is provided in Algorithm 1 of the Appendix.
2.3.3 Transformer-Block-Level Decompression
We now have a complete recipe for decompressing entropy-coded exponents in a massively parallel manner. During inference, the LLM weights stored in DFloat11 format, along with auxiliary variables (the thread-level gap array and block-level output position array), reside entirely in GPU memory. When a weight matrix is needed for matrix multiplication, it is decompressed on-the-fly into the original BFloat16 format. Once the matrix multiplication is complete, the BFloat16 matrix is immediately discarded to conserve GPU memory.
In practice, decompressing a single weight matrix often underutilizes GPU resources due to its relatively small size. As the matrix size increases, decompression throughput improves. Figure 7 illustrates this trend, showing how DFloat11 decompression scales with matrix size. To capitalize on this, we propose batching the decompression of multiple matrices together to improve throughput and hide latency. Specifically, we decompress all DFloat11 weight matrices within a transformer block as a single batch. This batched decompression occurs right before the forward pass of the transformer block. We also compress the token embedding and language modeling head of LLMs. Since these matrices are large enough to saturate GPU resources, batching their decompression is unnecessary.
3 Experiments
Table 1: DF11 statistics for various models. Model sizes are shown before and after compression.
Model Original $\rightarrow$ DF11 Compressed Compression Ratio Avg. Bit Width
Large Language Models
Llama 3.1 8B Instruct 16.06 GB $\rightarrow$ 10.90 GB 67.84% 10.85
Llama 3.3 70B Instruct 141.11 GB $\rightarrow$ 95.40 GB 67.61% 10.82
Llama 3.1 405B Instruct 811.71 GB $\rightarrow$ 551.22 GB 67.91% 10.87
Qwen 3 14B 29.54 GB $\rightarrow$ 20.14 GB 68.17% 10.91
QwQ 32B 65.53 GB $\rightarrow$ 44.65 GB 68.14% 10.90
Mistral Nemo Instruct 24.50 GB $\rightarrow$ 16.59 GB 67.74% 10.84
Mistral Small 3 47.14 GB $\rightarrow$ 31.86 GB 67.58% 10.81
Phi 4 Reasoning Plus 29.32 GB $\rightarrow$ 19.83 GB 67.64% 10.82
DeepSeek R1 Distill Llama 8B 16.06 GB $\rightarrow$ 10.89 GB 67.81% 10.85
Diffusion Transformers
FLUX.1 dev 23.80 GB $\rightarrow$ 16.33 GB 68.61% 10.98
FLUX.1 schnell 23.78 GB $\rightarrow$ 16.31 GB 68.58% 10.97
Stable Diffusion 3.5 Large 16.29 GB $\rightarrow$ 11.33 GB 69.52% 11.12
We empirically evaluate the effectiveness of DF11 compression and its GPU inference efficiency. A range of recent LLMs and DMs are compressed from their original BFloat16 format into DF11, and we report the resulting compression ratios. We then compare the inference performance of DF11-compressed models against their uncompressed counterparts across different GPUs, followed by an ablation study to analyze the impact of compression.
Software and Hardware We implement the DF11 decompression kernel in CUDA and C++, and integrate it into the HuggingFace Transformers huggingface inference framework. We evaluate the inference efficiency of our DF11 models against the original BF16 counterparts. We use the HuggingFace Accelerate framework to support CPU offloading and multi-GPU inference. To assess the performance of the DF11 kernel across different hardware configurations, we run experiments on multiple machines with varying GPU and CPU setups. The hardware specifications for all experimental machines are provided in Table 5 in the Appendix.
Table 2: Comparison of accuracy and perplexity for the BF16 and DF11 models on different benchmarks. DF11 compression results in absolutely no loss in accuracy or perplexity.
Accuracy Perplexity
Model Data Type MMLU TruthfulQA WikiText C4
Llama 3.1 8B Instruct BF16 68.010 $\pm$ 0.375 36.965 $\pm$ 1.690 8.649 21.677
DF11 (Ours) 68.010 $\pm$ 0.375 36.965 $\pm$ 1.690 8.649 21.677
Table 3: Comparison of peak GPU memory usage and text-to-image generation time for diffusion transformers in BF16 and DF11, using a single A5000 GPU.
Peak GPU Memory (GB) Generation Time (s)
Model BF16 DF11 (Ours) BF16 DF11 (Ours)
Stable Diffusion 3.5 Large 16.44 11.78 66.36 $\pm$ 0.13 69.08 $\pm$ 0.11
FLUX.1 dev 23.15 16.72 74.41 $\pm$ 0.15 78.53 $\pm$ 0.18
3.1 Results
Figure 4: Comparison of throughput (left three) and latency (right three) for token decoding using the original BF16 models and their DF11-compressed counterparts. Portions of the BF16 models are offloaded to the CPU due to GPU memory constraints.
[…]
DF11 compresses models to 70% size. Table 1 presents the compression factors of DF11 for a wide selection of recent LLMs and DMs. Specifically, we apply compression to all weight matrices and token embeddings in LLMs and all weight matrices in the transformer blocks of DMs. The models we compress include Llama 3.1/3.3 grattafiori2024llama , Qwen 3 yang2024qwen2 , Mistral Nemo/Small mistral2024nemo ; mistral2025small3 , Phi 4 phi4 , DeepSeek R1 Distilled guo2025deepseek , Stable Diffusion 3.5 stablediffusion35 , FLUX.1 flux . DF11 achieves approximately 70% compression across all models, corresponding to an effective bit width of around 11 bits.
Accuracy and perplexity evaluations confirm DF11 compression is lossless. We verify the lossless property of DF11 compression through a series of accuracy and perplexity evaluations on standard benchmarks. Evaluations are conducted using lm_evaluation_harness eval-harness , reporting accuracy on MMLU mmlu and TruthfulQA truthfulqa , and word-level perplexity on WikiText wikitext and C4 c4 . The results are shown in Table 2. As demonstrated, the compressed model achieves identical accuracy and perplexity to the original BF16 counterpart. We also present the text-to-image generation results of BF16 and DF11 Stable Diffusion 3.5 Large model in Appendix J. Given the same random seed and text prompt, the image generated are pixel-wise identical with the original model.
[…]
DF11 reduces memory usage for diffusion transformers with minimal latency impact. We evaluate the effect of DF11 compression on diffusion transformer models by measuring peak GPU memory and text-to-image generation time for an $\times1024$ image, over 5 runs. As shown in Table 3, DF11 reduces memory consumption by 28.3% for Stable Diffusion 3.5 and 27.8% for FLUX.1. The relative increase in latency is small: 4.1% for Stable Diffusion and 5.5% for FLUX.1.
[…]
Figure 6: Comparison of latency breakdown for DF11 and BF16 Llama 3.1 8B Instruct during GPU inference for different token batch sizes, using one A100-40GB GPU.
Figure 7: Throughput (left two) and latency (right two) comparisons between transferring BF16 matrices from CPU to GPU and decompressing the same matrices on GPU using the NVIDIA nvCOMP ANS library and our proposed DF11 kernel, across matrix sizes and GPU types.
3.2 Ablation Study
Latency breakdown shows decompression overhead is amortized at larger batch sizes. We analyze the latency of Llama 3.1 8B Instruct in BF16 and DF11 formats across varying token batch sizes on an A100-40GB GPU. For each setting, we measure the average latency of each component over 10 runs, as shown in Figure 6. DF11 introduces additional latency from decompressing the token embedding, transformer blocks, and language modeling head. This overhead is constant and independent of batch size, so increasing the token batch size effectively amortizes the cost.
[…]
4 Related Works
Data Formats for Model Weights Full-precision model weights are typically stored in formats such as BF16, FP16, or FP32. Several works have proposed 4-bit compressed formats, including FP4, INT4, NF4 (NormalFloat) dettmers2023qlora , AF4 (AbnormalFloat) yoshida2023af4 , and SF4 (Student Float) dotzel2024learning , which represent each parameter with 4 bits. Unlike these lossy formats, the proposed DF11 format compresses weights losslessly.
Lossless Model Compression While lossy compression methods such as pruning frantar2023sparsegpt and quantization lin2024awq ; frantar2022gptq are well-studied, lossless compression remains less explored. Four prior works have addressed this area. Deep Compression han2015deep_comp applied Huffman coding huffman_coding to quantized CNNs, achieving 22% additional compression. ZipNN hershcovitch2024zipnn extended this approach to language models with improved compression over classical methods. However, both techniques target storage efficiency and do not support inference-time gains. NeuZip hao2024neuzip is the only prior work supporting GPU inference. It uses Asymmetric Numeral Systems (ANS) with layer-wise decompression and relies on NVIDIA’s nvCOMP for GPU-based operations. nvCOMP is no longer open source, and its binary-only distribution limits adoption. Moreover, as shown in Figure 7, nvCOMP ANS incurs higher latency and lower throughput compared to our DFloat11 kernel. Huff-LLM yubeaton2025huffllm is designed for FPGA-like hardware and is not applicable to GPUs. Additional discussion of related formats is presented in Appendix B.
5 Conclusion
We present Dynamic-Length Float (DFloat11), a lossless compression framework for efficient GPU inference of BFloat16 models. DFloat11 leverages the information inefficiency of foundation model weights through entropy-coded, dynamic-length encoding, achieving compression near the information-theoretic limit. We develop hardware-aware algorithms to support fast inference with compressed weights. Extensive evaluations show that DFloat11 lowers GPU memory requirement for efficiently running LLMs and DMs, enables longer generation lengths, while introducing minimal decompression overhead and preserving bit-exact accuracy.
Appendix
Appendix A Discussion: Is Quantization a Universal Solution?